# Interpreting Fedspeak with Confidence: A LLM-Based Uncertainty-Aware Framework Guided by Monetary Policy Transmission Paths

### v2

@@ Introduction
Traditional methods in financial sentiment analysis often face a trade-off between performance and interpretability. For instance, dictionary-based approaches are simple and highly interpretable, but their performance is often limited as they struggle to understand complex contexts or incorporate external knowledge. In contrast, language models like FinBERT (Araci 2019; Liu et al. 2021), after being fine-tuned with domain-specific supervised data (Shah, Paturi, and Chava 2023), can capture contextual nuances effectively and achieve superior performance. However, the black-box nature of such model leads to a lack of transparency in their decision-making process, resulting in poor interpretability.
@@
Therefore, to align the model behavior with that of human analysts, we augment the input texts by extracting financial entity relations and reasoning over monetary policy transmission paths using structured templates. This augmentation is designed both to emulate expert-level economic reasoning and to reduce the model’s cognitive risk by compensating for missing domain-specific knowledge.
Then, we redefine the model prediction uncertainty as perceptual uncertainty (PU), decomposed into cognitive risk (CR) and environmental ambiguity (EA). Building on this formulation, we introduce a dynamic uncertainty decoding module that adapts model behavior based on the PU level. We tune PU–related hyperparameters on the validation set and use them to evaluate the model prediction confidence.
@@
Recent applications of LLMs in the financial domain have led to the emergence of domain-specific models. proprietary models like BloombergGPT (Wu et al. 2023) and open-source models such as FinGPT (Liu et al. 2023) are often instruction-tuned to perform financial tasks such as sentiment analysis. Existing financial evaluation benchmarks, such as FinQA (Chen et al. 2021), FinBen (Xie et al. 2024), FinTextQA (Chen et al. 2024), and FinDER (Choi et al. 2025), are designed to assess the performance of LLMs on a diverse range of tasks. These tasks cover numerical reasoning, retrieval-augmented generation evaluation, general understanding, and financial regulation and compliance. These benchmarks establish a standard for measuring the reliability of LLMs in various financial tasks.
@@
Fedspeak plays a pivotal role in shaping market expectations and asset prices, offering key insights into the Federal Reserve’s economic outlook and policy intentions (Ehrmann and Talmi 2020; Cieslak and Vissing-Jorgensen 2021; Mathur et al. 2022; Gómez-Cram and Grotteria 2022). Its nuanced language requires careful interpretation, as tone and sentiment in communications can significantly affect financial variables and market volatility (Gorodnichenko, Pham, and Talavera 2023; Curti and Kazinnik 2023). Approaches to analyzing Fedspeak have evolved from early dictionary-based methods (Loughran and McDonald 2011) and topic modeling techniques (Boukus and Rosenberg 2006; Edison and Carcel 2021) to Transformer-based models that better capture contextual subtleties (Liu et al. 2023; Shah, Paturi, and Chava 2023). LLM-based methods have become increasingly popular for policy stance analysis (Peskoff et al. 2023; Hansen and Kazinnik 2024). Domain-specific adaptations have shown promise in improving classification accuracy (Gambacorta et al. 2024; Geiger et al. 2025). However, most existing approaches limit LLMs to predicting stance labels, underutilizing their potential.
@@
When setting monetary policy, the Federal Reserve relies on deliberations and votes by FOMC members to reach collective decisions. These decisions reflect an aggregation of individual views and preferences. Committee members hold different policy stances: Members who prefer supporting increased output and employment and exhibit a stronger aversion to unemployment risk are considered dovish. Members who prefer controlling inflation and exhibit stronger aversion to inflation risk are considered hawkish (Bordo and Istrefi 2023). From a trade-off perspective, dovish members place more weight on maximizing employment, whereas hawkish members assign more weight to maintaining price stability (Shah, Paturi, and Chava 2023; Smales 2023). We follow prior literature in assigning stance labels based on these definitions. We define policy stances as dovish, hawkish, and neutral. The detailed definitions and explanations are illustrated in the detailed data augmentation section of the appendix.
@@
Existing research on monetary policy primarily focuses on understanding how policy tools, such as interest rate adjustments, affect key economic targets like inflation and employment. These studies trace the monetary policy transmission paths through the financial system and the real economy, offering a theoretical foundation for policy design (Passos, Carrasco-Gutierrez, and Loureiro 2024; Bosshardt et al. 2024; Suh 2025; Alessandri, Òscar Jordà, and Venditti 2025). Given the importance of monetary policy transmission paths, we construct the transmission path: Define a quadruple $\mathbf{\Gamma}=(\mathbf{X},\mathbf{Y},\mathbf{Z},\mathbf{M})$, where $\mathbf{X}\in{R}^{n}$ is an $n$ dimensional shock vector representing economic phenomena, policy implementation, policy framework shifts, or external shocks, serving as the starting point; $\mathbf{Y}=\{y_{1},y_{2},\dots,y_{k}\}$ is the set of monetary policy transmission channels, with $k$ being the number of channels, including the credit channel, asset price channel, aggregate demand channel, etc.; $\mathbf{Z}=\{Z_{1},Z_{2},\dots,Z_{p}\}$ is the monetary policy transmission path, defined as the sequence of expectation shifts or market/economic indicator responses within a specific transmission channel. For each channel $y_{i}\in\mathbf{Y}$, there exists a transmission path $Z_{y_{\text{i}}}$, which consists of a series of state transitions, as shown in Equation 3:
@@
LLMs may generate unreliable information. During next-token prediction, reliability can be affected by ambiguity in the input data and insufficient domain knowledge. We treat our model as a policy analyst. Policy analysts face reliability issues in decision-making due to EA and CR. EA stems from data-related uncertainty, such as distributional shifts, semantic ambiguity, and other stochastic variations in the input (Walters et al. 2023; le Roux and Bopp 2025). EA causes policy analysts to consider multiple alternative options during decision-making. CR arises from limited domain knowledge, missing information, or skill gaps (Walters et al. 2023). CR causes policy analysts to make unreliable judgments due to cognitive limitations. Higher levels of EA or CR increase the policy analyst’s PU. We incorporate a measure of PU into our model framework. Let the LLM be denoted as $\mathcal{M}$, and let the prompt be tokenized into vector representations $\mathbf{q}$. We use the top-$k$ logits from $\mathcal{M}$ to construct a Dirichlet distribution (Sensoy, Kaplan, and Kandemir 2018; Ma et al. 2025), as shown below:
@@
Table 1: Performance comparison of our method against zero-shot and fine-tuned baselines on the FOMC dataset.
@@
We utilized the open-source ModelScope Swift (Zhao et al. 2025) framework for all model training and inference. All base models were fine-tuned using the LoRA method. During inference, we applied greedy decoding and incorporated FlashAttention for efficiency. For dynamic uncertainty decoding, hyperparameters were selected based on validation set performance. The search space was defined as follows: Top-K: {3, 10, 15, 20, 25, 30}; Threshold percentiles: {1, 0.95, 0.9, 0.85, 0.8, 0.75, 0.7}; Sampling temperature: {0.1, 0.2, 0.3, 0.4, 0.5, 1.0, 1.5, 2.0}. All stochastic components were controlled with a fixed random seed 42 to ensure reproducibility. All experiments were conducted on four NVIDIA A800 80GB GPUs. Full implementation details are provided in the detailed experiment setup section of the appendix.
@@
We train and evaluate our method on the FOMC (Shah, Paturi, and Chava 2023) dataset, which contains three distinct categories of Federal Reserve communications: meeting minutes, press conference transcripts, and speeches. The dataset spans from January 1996 to October 2022, covering multiple economic cycles, including the dot-com bubble, the 2008 financial crisis, and the COVID-19 pandemic period. We report two metrics for evaluations and comparisons: Macro-F1 and Weighted-F1. Weighted-F1 is the original metric established for the FOMC dataset, ensuring consistency with previous research and enabling fair comparison with existing baselines. It accounts for the varying sample sizes across different classes by weighting each class’s F1 score according to its support. The Weighted-F1 score is calculated as:
$\text{Weighted-F1}=\frac{\sum_{i=1}^{C}(s_{i}\cdot F1_{i})}{\sum_{i=1}^{C}s_{i}}$ (11)
where $C$ is the number of classes, $s_{i}$ is the support (number of true instances) for class $i$, and $F1_{i}$ is the F1 score for class $i$. Specifically, the F1 score for each class is computed as:
$F1_{i}=\frac{2\cdot\text{Precision}_{i}\cdot\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (12)
We also report Macro-F1 to provide a complementary view of model performance. While Weighted-F1 emphasizes performance on majority classes, Macro-F1 treats all classes equally regardless of their frequency in the dataset. The Macro-F1 score is calculated as:
$\text{Macro-F1}=\frac{1}{C}\sum_{i=1}^{C}F1_{i}$ (13)
where $C$ is the number of classes and $F1_{i}$ is the F1 score for class $i$.
@@
We compare our method with several widely used language models (Achiam et al. 2023; Abdin et al. 2024; GLM et al. 2024; Yang et al. 2025; Guo et al. 2025; Comanici et al. 2025) and previous approaches (Liu et al. 2021; Peskoff et al. 2023; Fanta and Horvath 2024). For open-sourced language models, we report both zero-shot and fine-tuned results respectively. The results presented in Table 1 demonstrate the effectiveness of our proposed approach across all three categories of Federal Reserve communications. Our proposed method achieves substantial improvements over all baselines, demonstrating the effectiveness of our approach for financial sentiment analysis. Specifically, our method achieves 0.7327 Macro-F1 and 0.7426 Weighted-F1 on the combined dataset, representing significant improvements of 6.6% and 6.2% respectively over the strongest baseline (0.6662 and 0.6802). These results highlight the effectiveness of our methods.
@@ Conclusion
We introduce a domain-specific reasoning approach grounded in the monetary policy transmission mechanism to emulate how human experts analyze policy stances. We propose a dynamic uncertainty decoding module that effectively identifies cases where the model lacks sufficient knowledge or exhibits contextual confusion, rendering it unable to provide viable predictions. Under high perceptual uncertainty, our approach enables the model to identify unreliable outputs, improving reliability under realistic decision-making conditions. We validate our approach through a comprehensive hyperparameter search and demonstrate that the framework achieves state-of-the-art performance on the policy stance analysis task. Beyond improved robustness, our framework also supports human analysts by highlighting predictions that are well-grounded versus those requiring caution. Our work offers a new direction for evaluating the reliability, transparency, and economic interpretability of LLM predictions in policy stance analysis.
@@
While our approach demonstrates that incorporating domain-specific knowledge and dynamic decoding significantly improves model performance and reliability, it still relies on carefully designed domain templates and faces limitations in runtime efficiency and strategy design. Additionally, the current framework requires further enhancement to better handle samples involving contextual confusion or implicit statements. We leave these challenges for future work, including improvements in reasoning accuracy and framework scalability.

### v3

@@ Introduction
Traditional methods in financial sentiment analysis often face a trade-off between performance and interpretability. For instance, dictionary-based approaches are simple and highly interpretable, but their performance is often limited as they struggle to understand complex contexts or incorporate external knowledge. In contrast, language models like FinBERT (Araci 2019; Liu et al. 2021), after being fine-tuned with domain-specific supervised data (Shah et al. 2023), can capture contextual nuances effectively and achieve superior performance. However, the black-box nature of such model leads to a lack of transparency in their decision-making process, resulting in poor interpretability.
@@
Therefore, to align the model behavior with that of human analysts, we augment the input texts by extracting financial entity relations and reasoning over monetary policy transmission paths using structured templates. This augmentation is designed both to emulate expert-level economic reasoning and to reduce the model’s cognitive risk by compensating for missing domain-specific knowledge. Then, we redefine the model prediction uncertainty as perceptual uncertainty (PU), decomposed into cognitive risk (CR) and environmental ambiguity (EA). Building on this formulation, we introduce a dynamic uncertainty decoding module that adapts model behavior based on the PU level. We tune PU–related hyperparameters on the validation set and use them to evaluate the model prediction confidence.
@@
Recent applications of LLMs in the financial domain have led to the emergence of domain-specific models. proprietary models like BloombergGPT (Wu et al. 2023) and open-source models such as FinGPT (Liu et al. 2023) are often instruction-tuned to perform financial tasks such as sentiment analysis. Existing financial evaluation benchmarks, such as FinQA (Chen et al. 2021), FinBen (Xie et al. 2024), FinTextQA (Chen et al. 2024a), and FinDER (Choi et al. 2025), are designed to assess the performance of LLMs on a diverse range of tasks. These tasks cover numerical reasoning, retrieval-augmented generation evaluation, general understanding, and financial regulation and compliance. These benchmarks establish a standard for measuring the reliability of LLMs in various financial tasks. Another line of research enhances LLM reasoning by decomposing complex and ambiguous tasks into structured, domain-informed components, as seen in FinEntity, EFSA, and DEFINE (Tang et al. 2023; Chen et al. 2024b; Hu et al. 2024). These works highlight the value of integrating domain priors with structured task design to improve accuracy and reliability.
@@
Fedspeak plays a pivotal role in shaping market expectations and asset prices, offering key insights into the Federal Reserve’s economic outlook and policy intentions (Ehrmann and Talmi 2020; Cieslak and Vissing-Jorgensen 2021; Mathur et al. 2022; Gómez-Cram and Grotteria 2022). Its nuanced language requires careful interpretation, as tone and sentiment in communications can significantly affect financial variables and market volatility (Gorodnichenko et al. 2023; Curti and Kazinnik 2023). Approaches to analyzing Fedspeak have evolved from early dictionary-based methods (Loughran and McDonald 2011) and topic modeling techniques (Boukus and Rosenberg 2006; Edison and Carcel 2021) to Transformer-based models that better capture contextual subtleties (Liu et al. 2023; Shah et al. 2023). LLM-based methods have become increasingly popular for policy stance analysis (Peskoff et al. 2023; Hansen and Kazinnik 2024). Domain-specific adaptations have shown promise in improving classification accuracy (Gambacorta et al. 2024; Geiger et al. 2025). However, most existing approaches limit LLMs to predicting stance labels, underutilizing their potential.
@@
When setting monetary policy, the Federal Reserve relies on deliberations and votes by FOMC members to reach collective decisions. These decisions reflect an aggregation of individual views and preferences. Committee members hold different policy stances: Members who prefer supporting increased output and employment and exhibit a stronger aversion to unemployment risk are considered dovish. Members who prefer controlling inflation and exhibit stronger aversion to inflation risk are considered hawkish (Bordo and Istrefi 2023). From a trade-off perspective, dovish members place more weight on maximizing employment, whereas hawkish members assign more weight to maintaining price stability (Shah et al. 2023; Smales 2023). We follow prior literature in assigning stance labels based on these definitions. We define policy stances as dovish, hawkish, and neutral. The detailed definitions and explanations are illustrated in the detailed data augmentation section of the appendix.
@@
Existing research on monetary policy primarily focuses on understanding how policy tools, such as interest rate adjustments, affect key economic targets like inflation and employment. These studies trace the monetary policy transmission paths through the financial system and the real economy, offering a theoretical foundation for policy design (Passos et al. 2024; Bosshardt et al. 2024; Suh 2025; Alessandri et al. 2025).
Given the importance of monetary policy transmission paths, we construct the transmission path: Define a quadruple $\mathbf{\Gamma}=(\mathbf{X},\mathbf{Y},\mathbf{Z},\mathbf{M})$, where $\mathbf{X}\in{R}^{n}$ is an $n$ dimensional shock vector representing economic phenomena, policy implementation, policy framework shifts, or external shocks, serving as the starting point; $\mathbf{Y}=\{y_{1},y_{2},\dots,y_{k}\}$ is the set of monetary policy transmission channels, with $k$ being the number of channels, including the credit channel, asset price channel, aggregate demand channel, etc.; $\mathbf{Z}=\{Z_{1},Z_{2},\dots,Z_{p}\}$ is the monetary policy transmission path, defined as the sequence of expectation shifts or market/economic indicator responses within a specific transmission channel. For each channel $y_{i}\in\mathbf{Y}$, there exists a transmission path $Z_{y_{\text{i}}}$, which consists of a series of state transitions, as shown in Equation 3:
@@
LLMs may generate unreliable information. During next-token prediction, reliability can be affected by ambiguity in the input data and insufficient domain knowledge. We treat our model as a policy analyst. Policy analysts face reliability issues in decision-making due to EA and CR. EA stems from data-related uncertainty, such as distributional shifts, semantic ambiguity, and other stochastic variations in the input (Walters et al. 2023; le Roux and Bopp 2025). EA causes policy analysts to consider multiple alternative options during decision-making. CR arises from limited domain knowledge, missing information, or skill gaps (Walters et al. 2023). CR causes policy analysts to make unreliable judgments due to cognitive limitations. Higher levels of EA or CR increase the policy analyst’s PU. We incorporate a measure of PU into our model framework. Let the LLM be denoted as $\mathcal{M}$, and let the prompt be tokenized into vector representations $\mathbf{q}$. We use the top-$k$ logits from $\mathcal{M}$ to construct a Dirichlet distribution (Sensoy et al. 2018; Ma et al. 2025), as shown below:
@@
Table 1: Performance comparison of our method against zero-shot and fine-tuned baselines on the FOMC dataset. Zero-Shot denotes the baseline models; Fine-tuned applies LoRA fine-tuning on FOMC data to the Zero-Shot models; Ours further integrates the proposed data augmentation and uncertainty decoding module with the Fine-tuned Qwen-3-14B.
Experiment
@@
We utilized the open-source ModelScope Swift (Zhao et al. 2025) framework for all model training and inference. All base models were fine-tuned using the LoRA method. During inference, we applied greedy decoding and incorporated FlashAttention for efficiency. For dynamic uncertainty decoding, hyperparameters were selected based on validation set performance. The search space was defined as follows: Top-K: {3, 10, 15, 20, 25, 30}; Threshold percentiles: {1, 0.95, 0.9, 0.85, 0.8, 0.75, 0.7}; Sampling temperature: {0.1, 0.2, 0.3, 0.4, 0.5, 1.0, 1.5, 2.0}. All stochastic components were controlled with a fixed random seed 42 to ensure reproducibility. All experiments were conducted on four NVIDIA A800 80GB GPUs. Full implementation details are provided in the appendix.
@@
We train and evaluate our method on the trillion dollar Words  (Shah et al. 2023) FOMC dataset, which contains three distinct categories of Federal Reserve communications: meeting minutes, press conference transcripts, and speeches. The dataset spans from January 1996 to October 2022, covering multiple economic cycles, including the dot-com bubble, the 2008 financial crisis, and the COVID-19 pandemic period. We report two metrics for evaluations and comparisons: Macro-F1 and Weighted-F1. Weighted-F1 is the original metric established for the FOMC dataset, ensuring consistency with previous research and enabling fair comparison with existing baselines.
@@
We compare our method with several widely used language models like GPT (Achiam et al. 2023), Phi (Abdin et al. 2024), GLM (Zeng et al. 2024), Qwen (Yang et al. 2025), Deepseek (Guo et al. 2025), and Gemini (Comanici et al. 2025). The comparison also includes previous approaches like HD-Dissent (Peskoff et al. 2023), AICBC (Fanta and Horvath 2024), and FinBERT (Liu et al. 2021). For open-sourced language models, we report both zero-shot and fine-tuned results respectively. The results presented in Table 1 demonstrate the effectiveness of our proposed approach across all three categories of Federal Reserve communications. Our proposed method achieves substantial improvements over all baselines, demonstrating the effectiveness of our approach for financial sentiment analysis. Specifically, our method achieves 0.7327 Macro-F1 and 0.7426 Weighted-F1 on the combined dataset, representing significant improvements of 6.6% and 6.2% respectively over the strongest baseline (0.6662 and 0.6802). These results highlight the effectiveness of our methods.
@@ Conclusion
We present a domain-specific reasoning framework grounded in the monetary policy transmission mechanism to emulate expert analysis of policy stances. We propose a dynamic uncertainty decoding module that effectively identifies cases where the model lacks sufficient knowledge or exhibits contextual confusion, rendering it unable to provide viable predictions. Under high perceptual uncertainty, our approach enables the model to identify unreliable outputs, improving reliability under realistic decision-making conditions. We validate our approach through a comprehensive hyperparameter search and demonstrate that the framework achieves state-of-the-art performance on the policy stance analysis task. Beyond improved robustness, our framework also supports human analysts by highlighting predictions that are well-grounded versus those requiring caution. Our work offers a new direction for evaluating the reliability, transparency, and economic interpretability of LLM predictions in policy stance analysis.
@@
While incorporating domain knowledge and dynamic decoding enhances performance and reliability, our approach still relies on hand-crafted templates and has runtime and strategy constraints. The main limitation is the limited availability of fine-grained data from central banks such as the BoE and ECB. Additionally, the framework needs improvement to better handle context ambiguity and implicit statements. We leave these challenges for future work.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62406267) and the Guangzhou Municipal Science and Technology Project (No. 2025A04J4070). Zhang acknowledges financial support from the National Science Foundation of China (No. 72373097).
@@ Appendix A Appendix
Data Augmentation Ablation
The ablation study is conducted with Gemini-2.5-Pro under an end-to-end evaluation setup.
Setting Weighted-F1
Full augmentation prompt 0.6695
w/o rules 0.5073
w/o examples 0.5024
w/o both 0.4950
Table 7: Ablation results for the augmentation prompt. Truncating or simplifying the template consistently reduces accuracy.
The ablation results show that both rules and examples play essential roles in the augmentation prompt. Removing either component leads to a substantial drop in Weighted-F1, and removing both results in the lowest performance. The full prompt provides the strongest guidance and produces the most reliable augmented data, confirming that the structured template is necessary for stable model behavior.